OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the second equation. 3x + y + 2z = 1 x + 2y – z = –8 3x – y – 3z = 18

OpenStudy (anonymous):

A.)7y – z = 25 5y – 6z = 42 B.) –5x – 5z = –10 6x – z = 19 C.)7x + 3z = –10 –5z = 19 D.) –5y + 5z = 25 –7y = 42

OpenStudy (anonymous):

so do you know how to isolate x in the second equation?

OpenStudy (anonymous):

@kbreezy102?